Find the height in inches of the free surface if 0.8 ft3 of water is poured into a conical tank 20 inches high with a base radius of 10 inches. How much additional water in in3 is required to fill the tank?

_______ inches above the tank Answer in 2 decimal places

_______ in3



1
Expert's answer
2021-08-17T16:12:25-0400

Volume of water (Vw) = 0.8 ft3 = 0.8 "\times" 1728 in3 = 1382.4 in3

Radius of cone (r) = 10 in

height of cone (h) = 20 in



Volume of larger cone

"V_l = \frac{1}{3} \pi r^2 h \\

= \frac{\pi \times 10^2 \times 20}{3} \\

= 2094.39 \;in^3"

Volume of smaller cone = "V_l - V_w"

"= 2094.39-1382.4 \\

= 711.99\; in^3"



from similar triangle concept

"\frac{r_0}{h_0}= \frac{10}{20} \\

r_0=2h_0"

Volume of smaller cone can be written as

"\frac{1}{3} \pi r^2_0h_0 = 711.99 \; in^3"

put the value of r o in the above equation

"\frac{1}{3} \pi (2h_0)^2h_0 = 711.99 \;in^3 \\

h_0 = 5.539 \;in"

thus the height of the free surface = "h- h_0"

"=20-5.539 \\

= 14.461 \;in"

and the additional water required to fill the cone will be equal to the volume of the smaller cone

"V_{addition} = 711.99 \;in^3"